OpenStudy (anonymous):

Write 26/10 as a whole or mixed number in simplest form.

OpenStudy (anonymous):

divide by 2 26/2 ------- 10/2 =12/5

OpenStudy (anonymous):

26/10 = 2.6 close whole number is 3 is that what you want

OpenStudy (anonymous):

wait now i get what you mean 12/5 = 2 2/5